### v1

# Eleanor Wiske Dillon Sonia Jaffe Microsoft Research Microsoft Research Nicole Immorlica Christopher T. Stanton Microsoft Research Harvard Business School, NBER, and CEPR

Abstract
We present evidence on how generative AI changes the work patterns of knowledge workers using data from a 6-month-long, cross-industry, randomized field experiment. Half of the 6,000 workers in the study received access to a generative AI tool integrated into the applications they already used for emails, document creation, and meetings. We find that access to the AI tool during the first year of its release primarily impacted behaviors that could be changed independently and not behaviors that required coordination to change: workers who used the tool spent 3 fewer hours, or 25% less time on email each week (intent to treat estimate is 1.4 hours) and seemed to complete documents moderately faster, but did not significantly change time spent in meetings.
Recent advances in generative AI have opened new possibilities for using AI to assist with or automate a variety of tasks. Early studies have already shown that these technologies can substantially increase workers’ productivity in targeted tasks (Brynjolfsson, Li and Raymond (2025), Peng et al. (2023), Noy and Zhang (2023), Dell’Acqua et al. (2023)). Over time, these advances have the potential to shift the nature of jobs (Eloundou et al., 2024) and perhaps the macroeconomy (Autor (2024), Acemoglu (2025)). However, the transition from the creation of new technologies to effective, widespread adoption in the workplace takes time (Brynjolfsson, Rock and Syverson, 2021). In lab experiments, participants are asked to focus on completing a single task and directed to use AI tools that are well-suited to that task. Outside the lab, workers must learn on their own which parts of their job benefit from the use of new tools and gradually change their habits of work. Larger productivity gains often require organizational changes to redesign processes and shift responsibilities across workers (Bresnahan, Brynjolfsson and Hitt, 2002).
In this paper, we present early evidence on how generative AI tools are changing the way knowledge workers do their jobs. Between September 2023 and October 2024, Microsoft partnered with 56 large firms to run a cross-industry field experiment to measure how access to an integrated generative AI tool, Microsoft 365 Copilot (hereafter, Copilot), changed work patterns. Each firm recruited workers for the study; workers were then randomly assigned to either receive access to this new generative AI tool or continue working with their current technologies. Participants in the study span many occupations, but all participating workers do some kind of work that involves heavy use of the Microsoft Office tools into which Copilot was integrated (e.g., emailing, video meetings, chat, document creation). Firms maintained the random assignment for six months, giving treated workers sufficient time to explore the new tool and integrate it into their daily work.
This study took place during the early rollout of Copilot, when firms were restricted to purchasing a small number of licenses [^1] Each company was able to purchase 300 licenses. In all cases, this represents only a small fraction of their total employees who used Microsoft’s Office suite.. This particular generative AI tool was new for all workers, and the modal worker in the study had no close coworkers with access to the tool. Firms varied in how much training in the new tool they provided to treated workers. These workers were then free to use or not use the tool during their regular work. Treated and control workers may have also used other generative AI tools during the study, for example ChatGPT had been publicly available for nearly a year by October 2023, but access to Copilot could make it easier for workers to use generative AI at work. Copilot integrates AI into applications these workers were already using regularly in their jobs, allowing some new uses, for example summarizing emails or searching for information within work content, and reducing the effort cost of using generative AI to create content within documents and emails.
We used highly detailed telemetry data collected by Microsoft products to track the behavior of the over 6,000 studied workers during the experiment and for several months before the introduction of the new tool. These telemetry measures capture the actions of workers, for example time spent reading emails or attending video meetings or time from creation to completion of documents. We do not observe the content of any work.
We use the product telemetry data to both confirm that workers assigned to the treatment group received a Copilot license and to track what share of workers used Copilot tools in any given week. Patterns of Copilot usage are interesting in themselves. The usage data also allow us to estimate treatment effects that adjust for low usage or noncompliance with the randomized assignment using an instrumental variable approach. We present these estimates of the local average treatment effect (LATE) alongside intent-to-treat estimates based only on workers’ assigned study arm.
On average, a worker assigned to receive Copilot used it at least once in 45% of the study weeks. Usage rates varied widely across firms, from a low of 21% of treated study workers using Copilot each week to a high of 75%. This variation is consistent with other surveys of AI adoption across firms (Bonney et al., 2024; McElheran et al., 2024), but the variation is notable given our sample is all large firms that were interested in being early adopters of this tool and participating in a study of its use. Industry and variations in workers’ roles (proxied by pre-study worker behavior) explain some of these differences. Nonetheless, indicators for a worker’s firm remain by far the most powerful predictor of usage, suggesting an important role for firm-specific differences that we are unable to measure such as training or managerial practices.
We find clear evidence that workers who use Copilot are beginning to change their work patterns in meaningful ways. During this early adoption phase, we see larger changes in behaviors that workers can adjust independently and less movement in behaviors that require coordination with colleagues.
Using Copilot regularly caused workers to spend nearly 3 hours less per week working on emails (1.4 hours less in the intent-to-treat estimates). This shift represents a 25% time savings from the 11.5 hours per week that the average studied worker spent on email in the pre-period. Copilot users also shifted when they worked on email, consolidating email work to open up two more hours of extended focus time each week. [^2] Appendix B.1 details how we define each of these outcomes. We find suggestive evidence that workers are completing the documents they author 6-20% faster. In contrast, while workers are most likely to use Copilot during video meetings, we find no changes for Copilot users in the time spent in meetings or the types of meetings attended. Similarly, we find only small and statistically insignificant increases in total time spent working in Word and the number of documents authored, indicating that workers have not shifted the nature of their work or taken on new responsibilities. We find few substantive differences in these patterns among workers with more close coworkers who also have access to Copilot or within firms with higher adoption rates, suggesting that larger shifts in responsibilities require time and broad institutional efforts, not just local team coordination.
This work joins a small set of early studies of generative AI in real workplaces, including Brynjolfsson, Li and Raymond (2025), Cui et al. (2024), Otis et al. (2023), and Toner-Rodgers (2024). These real-world studies complement a larger body of work understanding the impact of generative AI on controlled tasks in a lab setting. [^3] See, for example, Peng et al. (2023), Spatharioti et al. (2023), Vaithilingam, Zhang and Glassman (2022), Campero et al. (2022), Noy and Zhang (2023), Dell’Acqua et al. (2023), Cambon et al. (2023) These earlier studies all focus on the impact of AI tools on specific tasks or narrow categories of workers. Surveys indicate rapid adoption of generative AI at work across a much broader set of workers; Humlum and Vestergaard (2025) find 41% of Danish workers in exposed occupations used ChatGPT at work by the end of 2023 and Blandin, Bick and Deming (2024) find 24% of a representative sample of all U.S. workers used generative AI at work by August 2024. However, these surveys cannot measure the impact of generative AI use on work behavior. This study aims to bridge some of the gap between these two lines of research: measuring both take up and impacts of use for a generative AI tool on a broad set of workers, albeit at the cost of clear measures of worker productivity.
1 Patterns of Copilot Adoption
In this study, firms gave treated workers access to Copilot, but did not require them to use it in specific work tasks. [^4] Additional details on the experiment are in Appendix A. The median treated worker used Copilot in 42% of the weeks we followed them (the mean was 45%). [^5] Note that the data reflect engagement with Copilot, not with generative AI tools overall. These adoption rates may not be indicative of broader or long-run trends. During this early adoption period workers had less familiarity with the capacities of generative AI and the tools themselves were less developed, which may have lowered usage. However, firms may have recruited workers for this study who were particularly well suited to being early and enthusiastic adopters. Figure 1(a) shows the distribution of these individual usage rates. 2% of treated workers used Copilot every week, while 6% never used it at all. Figure 1(b) plots the share of treated workers using Copilot for each week since their firm allocated the experimental licenses. Usage trailed off modestly after a peak of 60% when workers were first granted access, presumably due to a combination of people wanting to try the new tool and the introductory training sessions offered by some firms. In the final two months, overall usage held steady at around 40%, for a mean usage rate of 44% of workers across the study.666 The mean weeks of Copilot use across workers differs slightly from the mean share of workers using Copilot in each event week because not all workers were followed for exactly six months. See Appendix A for more details on the experimental design.
Figure 1(b) also shows usage rates over time separately for Microsoft 365 Copilot usage in three programs: Teams, Outlook, and Word. Copilot can be accessed through Outlook, Teams, Word, Excel, PowerPoint, and as a separate M365 Chat, and the ”any” line in Figure 1(b) reflects total use across all apps. We focus our impact analysis on three applications because the early phase of prompt-based generative AI was most applicable to the tasks in these tools. Workers were most likely to use the Copilot in Teams, which can capture transcripts of video meetings and analyze them to generate meeting summaries and follow-up items or answer user queries like “what was discussed in the last 10 minutes?”. The next most common use was through Outlook, where Copilot can summarize email chains, review a worker’s inbox and flag emails that require a response, answer questions about information contained in emails more flexibly than a keyword search, and generate draft emails in response to prompts. The Word Copilot can draft text from prompts and answer questions about the contents of documents. This Copilot had the lowest usage of the three we study, partially because not all studied workers used Word regularly in their jobs (see Appendix B).
(a) Usage Rate by Individual
(b) Usage Rate by App
(c) Usage Rate by Industry
(d) Usage Rates across Apps by Firm Usage
Figure 1: Copilot Usage in the Treatment Group
Note: Figures (a) and (c) calculate the proportion of weeks in the post-period that each worker used Copilot and show the distribution across (a) workers overall or (c) workers by industry. In Figure (c), each solid dot represents 5 workers; lighter dots represent fewer than 5 workers. Mean industry usage rates are highlighted in red. Figure (b) shows average usage rates across all treated users over time. Figure (d) is a binscatter for the average share of Copilot users in a specific app in the average week grouped by the overall Copilot usage rate of the firm.
Some of the individual heterogeneity is also reflected in average firm usage rates, which ranged from as low as 21% to as high as 75%. [^7] Appendix Figure 3(a) shows how usage varied over time across firms. We do not observe many potentially important sources that likely drive this firm variation. For example, some firms likely provided training to workers on Copilot useage or may have altered managerial practices. However, we can look at how much of the variation in adoption is explained by (the proxies we have for) the type of work done at different firms (firm industry), the type of work done by study participants (pre-experiment individual work patterns [^8] Weekly averages of: emails read, unique email threads replied to, Teams meeting time, Teams meetings attended, time in Word, documents edited or read before completion, number of unique email recipients, number of unique people met 1-on-1.), and how firms distributed their Copilot licenses (the share of a worker’s coworkers who have a license). We cannot estimate the causal effect of any of these factors, only the extent to which they explain the observed variation.
Table 1: Predicting Adoption
(1) (2) (3) (4) (5) (6) (7)
Individ. pre-experiment $all\right>$ $all\right>$ $all\right>$ $all\right>$
behavior
Share of coworkers 0.044 0.053 0.099 0.038
with Copilot (0.024) (0.024) (0.023) (0.024)
[…]
Firm FEs Yes Yes Yes
Industry FEs Yes Yes
[…]
$R^{2}$ 0.093 0.094 0.235 0.189 0.184 0.050 0.139
Within $R^{2}$053 .006 .091
Note: This table reports OLS regressions where the dependent variable is the share of post-period weeks that an individual used Copilot. Coefficients for individual pre-experiment behavior variables are in Appendix Table A1.
Table 1 shows the $R$-squared from regressions of each treated worker’s average Copilot usage rate over the treatment period on various individual and firm metrics. The associated coefficients are in Appendix Table A1. Individual pre-experiment activity in Outlook, Teams, and Word explains only 9.3% of the variation. The share of a worker’s coworkers that have Copilot has a statistically significant impact on usage, but the effect is small and adds little explanatory power. Moving from no coworkers to the median of 12.5% would increase a person’s predicted usage by 0.55 percentage points. Firm fixed effects matter more, explaining 18.4% of the variation on their own and raising the $R$-squared from 0.094 to 0.235 when added to the other variables. Adoption rates vary meaningfully across industries (ranging from 30% in Construction and Manufacturing to 54% in Consumer Goods as shown in Figure 1(c)). Nevertheless, industry fixed effects explain less than a third as much of the variation in individual usage as firm fixed effects. [^9] Some of the firm variation could also come from time-of-year effects since firms joined the experiment at different times. In the Appendix we repeat this analysis with user-week level data where we can control for calendar time. The $R$-squared’s are overall much lower, since we are not averaging within a user, but firm fixed effects still add about twice as much explanatory power as individual pre-experiment work patterns.
Given the importance of the firm in determining Copilot usage, we also consider whether workers at firms with high uptake rates overall are also using Copilot differently. Figure 1(d) plots the firm average use intensity of Copilot within each of the three applications we study against the overall firm average Copilot usage. A relationship that parallels the 45°$\degree$ line, as is the case for Teams, indicates that this is an equally popular way to use Copilot at all firms. A flatter relationship, which we see in Outlook and Word, indicates that these applications are relatively more popular at firms with lower overall usage: the rate of Outlook and Word Copilot use rises more slowly than the increase in Copilot use overall. These patterns are consistent with our findings on the impacts of Copilot use. In Outlook and Word, Copilot can assist individual behaviors, so workers can find it useful even in environments where few colleagues are adopting this new tool. At firms with more active users, the additional usage seems to mainly come from the other apps. [^10] Appendix Figure 3(b) shows a similar pattern based on the average number of actions. As Copilot usage at a firm increases, the average number of Teams-Copilot actions among Copilot users increases, but the number of Outlook-Copilot actions is flat.
2 Effects of Generative AI on Patterns of Work
2.1 Methods
Since license assignment was randomized, comparing mean treatment and control group outcomes in the post-period would give an unbiased estimate. However, since our outcome metrics vary considerably across workers and over time, a difference-in-differences (DiD) approach that exploits our pre-Copilot observations greatly improves precision. Our intent-to-treat (OLS) specification estimates the effect of receiving a Copilot license on an outcome $Y_{it}$ for worker $i$ in week $t$ at firm $f$ using an event study design:
$Y_{it}=\alpha_{i}+\delta_{m_{t}}+\gamma_{f,\tau_{ft}}+\sum_{k\neq-1}\beta_{\tau}Z_{i}\cdot\mathbf{1}\{\tau_{ft}=k\}+\epsilon_{it}$ (1)
where $\delta_{m_{t}}$ captures calendar year-month effects, $\tau_{ft}$ is the event month for firm $f$ (month relative to when the firm started the experiment), and $Z_{i}$ is the worker’s experimental assignment. We include both calendar time and event time controls because firms entered and exited the study on a rolling basis. Note that while outcomes are measured on a weekly basis, we capture time-varying effects at a monthly frequency to reduce noise. We use Newey-West standard errors with a lag of 9 weeks to handle autocorrelation in outcomes.
Because of our experimental design, we can define time since “treatment” for every worker in the sample, both treated workers and their comparison control workers. The effects of Copilot are identified only by differences in the month-to-month within-worker changes in outcomes between the treated workers and the comparison workers at the same firm. The potential biases from a rolling DiD analysis highlighted by (Goodman-Bacon, 2021) and others are therefore not a concern in this setup. We also use a two-way fixed effect framework where a single $\beta$ on $Z_{i}\mathbf{1}(\tau_{ft}\geq 0)$ estimates the average treatment effect across all months of the post-period. This specification includes the same set of fixed effects, so the identifying variation is the same.
As highlighted in the previous section, many workers who had access to Copilot did not use it regularly, and some workers who were assigned to the control group nonetheless obtained Copilot licenses. To adjust for low usage and non-compliance we also follow Duflo (2001) and Hudson, Hull and Liebersohn (2017) and estimate an instrumented difference-in-difference specification. In this case, we use treatment assignment as an instrument for whether a worker was an above-median Copilot user, using Copilot in at least 42% of the post-treatment weeks. The single treatment effect regression equations are [^11] Note that $D_{it}$ does not vary (for an individual) across weeks within the pre- or post- period. It is one for all post-period weeks for individuals who were above-median Copilot users and zero otherwise. Because of the time fixed effects, $\hat{D}_{it}$ can vary month-to-month.
$Y_{it}$ $=\alpha_{i}+\delta_{m_{t}}+\gamma_{f,\tau_{ft}}+\beta\hat{D}_{it}+\epsilon_{it}$ (2)
$\hat{D}_{it}$ $=\tilde{\alpha}_{i}+\tilde{\delta}_{m_{t}}+\tilde{\gamma}_{f,\tau_{ft}}+\tilde{\beta}Z_{i}\cdot\mathbf{1}\{\tau_{ft}\geq 0\}+\eta_{it}.$
These estimates capture the local average treatment effect of Copilot for the complying workers who actually use it. To extend this IV setup to the multi-period event studies, we instrument each $\hat{D}_{it}$ with interactions of $Z_{i}$ and event time.
2.2 Average Effects
During this early adoption phase, the most prominent impact of Copilot on worker behavior was changes in how workers managed their email. In the pre-period, the average worker in our sample spent over 11 hours per week reading and responding to emails [^12] Details on the construction of all worker behavior measures are in Appendix B.1. , more than a quarter of a typical work week. For most workers, reading and writing emails does not represent their core job responsibilities but rather a communication overhead cost that supports their productive work. On average, workers with access to Copilot spent 1.4 hours less per week working on email in Outlook and workers who used Copilot regularly spent almost 3 fewer hours per week, a 25% decline from their pre-period average. Figure 2 illustrates that this reduction in Outlook time is apparent even in the first month after workers got access to the tool and persists throughout the study.
(a) Time in Outlook
(b) Time in Word
(c) Time in Teams Meetings
Figure 2: Event Study Plots for the Effect of Copilot on Time Allocation
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table 2.
Figure 2 also illustrates that we find no similar shifts in the time workers spend in Teams meetings or working in Word, despite the fact that these tasks are also complementary with the strengths of generative AI and, particularly in the case of Teams meetings, workers are regularly using Copilot through these applications. We can think of two hypothetical explanations for this difference. The first considers the demand for these tasks. When workers become more productive at a task, the effect on their total time spent doing it is ambiguous. Each unit of production in this task is now cheaper (in time) to produce, so workers may substitute towards it. For example, as generative AI improves workers’ ability to capture and summarize the output of meetings, workers might choose to do more work such as brainstorming or outlining documents live in a meeting rather than asynchronously. Alternatively, if there is no demand for more of this task, workers will complete the same work as before in less total time. For example, we might see meeting times shorten or recurring meetings repeat less frequently. It could be the case that the second effect dominates in the case of email management, but the two effects roughly balance out for video meetings and document writing.
The second explanation concerns coordination across teams and organizations. During this study, the treated workers had access to Copilot, but very few of their colleagues did. In 2023, almost no companies had developed, let alone implemented, systemic plans of how generative AI could transform their workflows and allocations of responsibilities across workers. While emailing involves communicating with others, the tasks of reading emails, writing emails, and tracking conversations that require follow-ups are entirely solitary. Each worker is free to develop her own process for managing her inbox. In contrast, meetings are group activities. Shifting what work gets done in meetings or how long they last requires coordinating with colleagues and agreeing on new norms. Writing one’s own contributions to a document is a solitary process, but taking on the responsibility of being the primary author of more documents or shifting work norms so that more documents are produced requires larger changes in organizational expectations.
Table 2: Effect of Copilot
Metric Pre-period mean (SD) OLS (SE) IV (SE) Workers $\mathbf{n}$
Outlook total session time 11.53 -1.35** -2.95** 5,803 281,283
(6.40) (0.069) (0.15)
Total concentration time 24.61 0.90** 2.05** 5,620 247,907
(6.19) (0.060) (0.14)
Outlook out-of-hours session time 2.08 -0.30** -0.65** 5,803 281,283
(1.65) (0.019) (0.041)
Unique email threads replied to 13.22 -0.23* -0.47* 6,116 325,472
(11.45) (0.081) (0.17)
Time to reply from email delivery 14.77 -0.31** -0.61** 6,111 285,332
(4.82) (0.092) (0.19)
[…]
Teams meeting time 5.21 0.041 0.081 5,698 208,754
(3.27) (0.034) (0.066)
Share of meetings attended late/early 0.30 0.006** 0.012** 5,645 178,218
(0.11) (0.002) (0.004)
Recurring Teams meetings time 2.32 0.013 0.025 5,698 254,319
(1.68) (0.016) (0.031)
[…]
Word total session time 1.35 0.018 0.032 2,150 114,863
(0.93) (0.023) (0.041)
Avg. document time to complete 175.56 -6.93 -11.65 2,150 36,523
(82.97) (5.37) (9.03)
Avg. collaborative document TTC 278.31 -35.83* -55.77* 1,673 11,033
(115.26) (13.99) (21.73)
Avg. non-collaborative document TTC 143.73 -3.72 -6.32 2,140 30,975
(77.58) (5.20) (8.83)
Completed documents 0.78 0.028 0.051 2,150 91,213
(0.75) (0.017) (0.031)
Completed collaborative documents 0.15 0.011 0.019 2,150 91,213
(0.22) (0.006) (0.010)
Significance: ** $q<0.01$, * $q<0.05$
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2) and the LATE coefficients (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$. All time metrics are in hours. Share of meetings attended late/early refers to the share of Teams meetings that a worker joined more than 5 minutes after the meeting started or left more than 5 minutes before it ended. We assign documents to a worker if they were the ‘primary editor’, making the more most edits to the document.
Looking in more detail at how workers changed their behavior when using Copilot, we find more evidence for the second hypothesis: workers were more likely to adjust behaviors that could be changed unilaterally and less likely to shift behaviors that required coordination with colleagues. In the top panel of Table 2, we see that workers’ responsiveness and intensity of email communication with colleagues is roughly unchanged. Workers who used Copilot replied to roughly the same number of conversations per week (13 vs. 13.5) and replied to emails slightly faster than before. These workers simply take less time to do roughly the same quantity of email work. To illustrate some of the benefits of this shift in behavior, note that workers who use Copilot not only spending less time in Outlook each week but also condense their work into fewer sessions. This opens up two more hours per week of concentration time, blocks of 30 minutes or more during the regular workday with no Outlook activity. Workers using Copilot also cut back on time spent in email outside of their regular work hours.
If the null effect on total time spent in meetings was masked by offsetting shifts in how workers were using meetings, we would expect to see some changes in the types of meetings workers attended after gaining access to Copilot. However, as shown in Table 2, we find no impact on time spent in recurring meetings relative to meetings overall. [^13] As shown in appendix table A3, we also find no significant effect on the number of meetings attended, ruling out a shift toward more shorter meetings, and on most meetings by type (long meetings with few participants, short meetings with many participants, etc.), which previous research has shown correlate with meeting purpose. While we do estimate small but statistically significant overall increases in large meetings for workers with Copilot access, the event studies in Figure A5 show that this increase is concentrated in the first month of Copilot access, which suggests these might be required training sessions in the new tool. The only impact of Copilot on meeting-related behavior that we find is a small, precisely estimated increase in the share of meetings where treated workers join late or leave early, an individual choice. Teams is the most common application in which workers used Copilot, so they likely found value in using generative AI in this context, perhaps in lowering effort to retain information from meetings or providing the ability to focus on participating instead of taking notes. However, this value did not translate into any visible shift in the structure of work during our sample.
Focusing on workers who were regular users of Word in the pre-period, the bottom panel of Table 2 provides more details on how Copilot access and use affects workers’ document production. We do not have the power to measure these effects very precisely because over half of study participants were not regular Word users. However, there are some suggestive patterns. Consistent with Noy and Zhang (2023), we find that workers with Copilot complete their documents more quickly, although the magnitude of the effect is far smaller than what they found in the lab, and we cannot reject the null hypothesis of no effect. Interestingly, the time savings are strongest, in both time saved (over 2 days) and percent reductions (20%), in documents where our studied worker was the primary editor but at least one other coworker also contributed. These multi-author documents take longer to complete, nearly 12 days on average, so there may be more scope for time savings. Workers may also feel more comfortable using generative AI staying on top of colleagues’ edits and outstanding tasks than drafting text in work documents. The larger effects in collaborative documents also reassure us that we are seeing some behavior shift beyond casual experimentation with a new tool. Consistent with workers changing the way they do their own work, but not taking on different responsibilities, we find no change in the number of documents where the studied workers are the primary editor.
2.3 Effects by Peer Copilot Use
To further explore the role of team and organizational coordination in shaping the early effects of generative AI, we consider whether the effects of Copilot access and use vary when workers have more close colleagues who also have access to Copilot. As discussed in Appendix B.2, we identify close colleagues as the coworkers our studied workers interact with most frequently through emails, meetings, and joint writing. Though we do not have random variation in the share of colleagues with Copilot, the treatment and control assignments are still random for people above (or below) the median, so we can get unbiased estimates for each subgroup.
Table 3: Effect of Copilot - Split by Share of Coworkers with Copilot licenses
Outcome Pre-period OLS (SE) IV (SE) $\mathbf{n}$
mean (SD)
Above Below Above Below Above Below Above Below
Outlook total session time 11.77 (5.98) 11.30 (6.78) -1.50** (0.095) -1.27** (0.10) -3.04** (0.19) -2.97** (0.23) 140,463 140,820
Total concentration time 24.53 (5.86) 24.69 (6.57) 1.05** (0.084) 0.78** (0.087) 2.22** (0.18) 1.90** (0.21) 128,272 119,635
Outlook out-of-hours session time 2.12 (1.55) 2.04 (1.75) -0.33** (0.026) -0.28** (0.027) -0.67** (0.053) -0.64** (0.062) 140,463 140,820
Unique email threads replied to 13.60 (11.43) 12.86 (11.38) -0.26 (0.11) -0.26 (0.12) -0.49 (0.21) -0.58 (0.26) 160,538 164,934
Time to reply from email delivery 15.00 (4.82) 14.54 (4.81) -0.25 (0.13) -0.35* (0.13) -0.47 (0.24) -0.75* (0.29) 144,189 141,143
[…]
Teams meeting time 5.71 (3.13) 4.72 (3.34) 0.009 (0.047) 0.067 (0.048) 0.017 (0.087) 0.14 (0.10) 105,376 103,378
% Teams meetings attended late/early 0.28 (0.096) 0.31 (0.12) 0.006 (0.003) 0.006 (0.003) 0.010 (0.004) 0.013 (0.006) 92,260 85,958
Recurring Teams meetings time 2.51 (1.62) 2.13 (1.74) -0.008 (0.022) 0.031 (0.022) -0.014 (0.041) 0.066 (0.046) 130,249 124,070
[…]
Word total session time 1.31 (0.91) 1.38 (0.95) 0.010 (0.031) 0.011 (0.033) 0.017 (0.055) 0.020 (0.061) 59,662 55,201
Avg. document time to complete 179.17 (86.74) 171.75 (77.25) -1.95 (7.69) -13.26 (7.71) -3.23 (12.74) -22.97 (13.35) 19,382 17,141
Avg. collaborative document TTC 280.76 (121.48) 275.15 (104.04) -36.29 (19.21) -24.95 (22.51) -54.44 (28.83) -42.13 (37.96) 6,507 4,526
Avg. non-collaborative document TTC 145.01 (77.50) 142.38 (77.72) 0.97 (7.48) -9.62 (7.46) 1.64 (12.63) -16.58 (12.85) 16,138 14,837
Completed documents 0.78 (0.75) 0.77 (0.75) -0.005 (0.023) 0.057 (0.025) -0.009 (0.041) 0.11 (0.046) 47,551 43,662
Completed collaborative documents 0.17 (0.23) 0.13 (0.21) 0.006 (0.008) 0.014 (0.007) 0.011 (0.014) 0.027 (0.014) 47,551 43,662
Significance: ** $q<0.01$, * $q<0.05$
Note: This table replicates the analysis in Table 2 separately for workers who had above or below median share of coworkers with Copilot licenses in the post-period. See that table not for explanations and variable definitions.
Table 3 splits our sample by whether our studied workers have above or below median share (12.5%) of close coworkers with access to Copilot. We see almost no meaningful differences in behavior between these two groups, with one exception in document creation. Workers who have fewer colleagues with access to Copilot author 14% more documents when they use Copilot themselves. These documents are primarily solo authored, since we see no increase in collaborative documents, and these solo documents are completed more quickly. This pattern suggests that some teams adapted in the short run and reallocated writing work to the one team member with access to the new generative AI tool. However, we do not expect this pattern to persist in the long run as generative AI tools become more widespread. Appendix Table A5 shows a similar exercise, splitting our sample into workers at firms with higher or lower overall Copilot adoption. Again, while there are some small differences in impacts across these two groups, we do not see evidence that firms with higher Copilot adoption showed more movement in behaviors that require group coordination. Overall, the lack of large differences suggests that the barriers to broader shifts in work patterns are not simply coordination costs within small teams but might include factors such as the time required to reimagine work flows and the institutional effort to drive these changes.
3 Conclusion
In order to capture the full productivity potential of new technologies, firms frequently have to undergo complex organizational transformations (Feigenbaum and Gross, 2024; Bresnahan, Brynjolfsson and Hitt, 2002). These institutional changes take time. Among recent technological innovations, generative AI is unique because it does not require new equipment or hardware to use, allowing individual workers to bring generative AI to work even without their employers’ support. As Blandin, Bick and Deming (2024) point out, the discrepancy between the high rates of generative AI use reported in surveys of workers, such as theirs, and the lower rates reported in surveys of firms, as in Bonney et al. (2024), suggests that workers use these tools even before their employers develop formal adoption strategies.
This study provides some of the earliest signals of how this individual use of generative AI is beginning to shape the patterns of work for knowledge workers. We observed workers during the first year in which generative AI tools were beginning to gain broad use at work. While their colleagues may have been using other generative AI tools, the workers we studied were often the only members of their team with access to Copilot during the early roll-out period. The changing behaviors we see are consistent with workers independently exploring these new tools and saving time on individual tasks, but the individual rollout did not exhibit the organization-wide shifts we might expect as teams and organizations begin larger transformations in response to this technology. We therefore expect that the patterns we find are only the beginning of a new era of work.
Appendix
Appendix A Experiment Details
More details on the study design and implementation are available in Dillon et al. (2025). At the time the experiment was run, firms could purchase at most 300 M365 Copilot licenses. The Microsoft marketing team recruited firms to join the experiment with the understanding that at least 50 of their licenses would be allocated at random between workers. Firms were asked to recruit at least twice the number of workers for the study as the number of Copilot licenses that would be randomly assigned. For example, a firm that chose to allocate 50 of their 300 licenses at random would need to recruit at least 100 workers to join the experiment, while a firm that chose to allocate 150 licenses at random would need to recruit 300 workers to join. Some firms had more than 400 workers participate in the study, while a few ended up with slightly less than 100. Each individual firm selected which workers they wanted to participate; some focused on a single department or work area while others pulled workers from many parts of the company. Firms were responsible for all communications with the participating workers, including obtaining necessary consent. Some firms provided additional information about the recruited workers such as department or job title. While the sample of workers with this additional information is too small for meaningful subgroup analyses, it confirms that the participating workers are generally early to mid-career knowledge workers such as analysts, project managers, or professional service providers.
Our analysis sample has 56 firms across a range of industries, as shown in Figure 1(a). There were 4 additional firms that agreed to participate in the study but did not follow through with randomization; these firms are excluded from the analysis sample. The experiment includes workers from many countries, with the majority based in the United States or Europe.
(a) Industry Mix
(b) Over Time
Figure A1: Participating Firms
Note: Figure (a) shows the number of participating firms in each industry. Figure (b) shows the number of firms active in the experiment over time.
For all firms, our data sample begins on May 5, 2023, providing at least 4 months of data before any Copilot licenses were allocated. Firms varied in when they allocated licenses, with the first starting in late September 2023, and the latest in April 2024. Figure 1(b) illustrates the number of firms actively maintaining randomization over the study window. We collected data on the workers at each firm for a maximum of six months after licenses were allocated. Six months is also the modal length of the post period across firms. A small number of firms have shorter treatment periods because they explicitly exited the experiment prior to the six-month post-period. In all analyses, we drop outcomes for the weeks ending December 29, 2023 and January 5, 2024 as work activity was low across all workers during this holiday period.
Appendix B Data Details
We analyzed millions of Microsoft 365 usage signals (telemetry), for both the treated and designated control workers, from product telemetry data that Microsoft collects. To maintain workers’ and firms’ privacy, we do not observe anything about the content produced by these workers, so our focus is on how workers allocate their time and interact with each other rather than on the quality of their work.
For some analyses, we restrict our estimation to a subsample of workers who used a particular application in question. While virtually every worker in our sample used Outlook regularly, fewer workers used Word, indicating that many workers likely do not have jobs that require document creation in Word. For results that examine document creation or interactions, we subset the analysis to workers with higher Word use in the pre-period, indicated by above-median time spent in Word (.27 hours/week) and number of documents originated as the ‘primary editor’ (.19 documents/week). For Teams meetings, we drop five firms that were not regular users of Teams prior to the experiment because they relied on a different video meeting program.
B.1 Outcome Metrics
We construct weekly metrics capturing each worker’s email, document editing, and meeting behaviors. Metrics are winsorized to the 95th percentile, except Word document count metrics, which are winsorized to the 99th percentile due to fewer weekly observations.
  • Outlook
    For email, we aggregate the signals of various ‘reading actions’ individuals take into ‘Outlook sessions’: blocks of time where they don’t go more than 15 minutes without reading an email.
    • Outlook total session time The time spent aggregated across all sessions in a week, measured in hours.
    • Outlook out-of-hours total session time The subset of Outlook session time that occurred outside the individual’s “working day” as defined as the 8 hours in which they were most active during one month of pre-period.
    • Total concentration time The total time spent during an individual’s working day in non-Outlook blocks at least 30 minutes long.
    • Unique email threads replied to For emails received that week that are replied to within 7 days, the number of unique email threads/conversations.
    • Time to reply from email delivery For emails received that week that are replied to within 7 days, the average length of time it takes to reply to an email.
    • Emails read Count of distinct emails read within a week.
    • Total emails in email threads replied to For emails received that week that are replied to within 7 days, the total number of emails in the threads/conversations.
  • Teams
    • Teams meeting total session time The time spent in scheduled Teams meetings (excluding spontaneous calls) in a week, measured in hours.
    • Share of meetings attended 5 min late/early The share of scheduled Teams meetings joined either $\geq$5 minutes after the Teams meeting was started or left $\geq$5 minutes before the meeting was ended.
    • Recurring Teams meetings total session time The time spent in Teams meetings marking as recurring meetings, measured in hours.
    • Total Teams meetings attended The number of Teams meetings attended in a week.
    • Total recurring meetings The number of Teams meetings that were marked recurring attended in a week.
    • Big short meetings The number of Teams meetings of length less than 1 hour with over 8 people.
    • Big long meetings The number of Teams meetings of length 1 hour or more with over 8 people.
    • Small short meetings The number of Teams meetings of length less than 1 hour with 8 or fewer people.
    • Small long meetings The number of Teams meetings of length 1 hour or more with 8 or fewer people.
  • Word
    • Across metrics, a worker is a primary editor if they make the most edits on a document; a document is completed when 90% of the total edits to the document have been completed. [^14] Some documents show stray edits long after most active work is done, which may be accidental. This 90% cutoff aligns most often with other indicators of completion such as attaching to an email or printing to pdf.
    • Word total session time The time spent in Word sessions, identified as blocks of actions with $<15$ minutes gap, measured in hours.
    • Avg. document time to complete The average time between the creation (first edit) and completion of a document, for documents completed that week and where the target worker is the primary editor, measured in hours.
    • Avg. collaborative document time to complete The same as the average time to complete, but only for documents with multiple collaborators.
    • Avg. non-collaborative document time to complete The same as the “avg. document time to complete” but only for documents with no collaborators (one editor).
    • Completed documents The number of documents completed in a week where the target worker is the primary editor.
    • Completed collaborative documents The same as “completed documents” but only for documents with multiple collaborators.
    • Others’ documents read/edited before completed The number of documents read or edited in a week for which the target worker is not the primary editor.
B.2 Coworkers
We also use the telemetry data to identify the closest coworkers of each person in the experiment, based on a weighted sum of their joint meetings, calls, emails, chats, and document collaborations. For each metric, we divide by the firm-specific average. We add these normalized numbers together for an interaction score and call someone a “coworker” if the score is greater than 8.0 (If they add the average amount of interaction, the score would be 5.0 since there are five metrics.) If an individual has more than 15 coworkers, we take only the 15 with the highest interaction score. The average worker has 9 coworkers. We can then calculate what share of an individual’s coworkers have a Copilot license. Figure A2 shows the distribution of that share.
Figure A2: Distribution of share of coworkers with Copilot
Appendix C Additional Tables and Figures
C.1 Adoption
Figure A3 shows some additional Copilot usage patterns. Figure 3(a) shows the variation across firms in usage over time. Figure A3 shows a binscatter of the mean weekly Copilot actions taken in each app among users using some form of Copilot that week by the firm-level share of users that are active in an average week. Similar to Figure 1(b), we see that the usage intensity grows with firm usage much more for Teams than for Outlook (with Word in between).
(a) Usage Rate by Firm
(b) Usage Intensity across Apps by Firm Usage
Figure A3: Copilot Usage
Note: Figure (a) shows usage over weeks of the experiment, with one line for each participating firm. Figure (b) shows a binscatter of the mean weekly Copilot actions taken in each app among users using some form of Copilot that week by the firm-level share of users that are active in an average week.
Table A1 presents all the coefficients associated with the regressions in Table 1. Table A2 does the same analysis with user-week data instead of user-level averages. The $R$-squared’s are lower since there is lots of unexplained week-to-week variation, however the pattern is similar.
Table A1: Predicting Adoption - Coefficients
Constant 0.3800∗∗ 0.3705∗∗
(0.0124) (0.0134)
Emails read -0.0003∗∗ -0.0003∗∗ -0.0002∗∗ -0.0004∗∗
($7.91\times 10^{-5}$) ($7.95\times 10^{-5}$) ($8.33\times 10^{-5}$) ($8.09\times 10^{-5}$)
Unique email threads replied to -0.0014 -0.0014 -0.0011 -0.0010
(0.0007) (0.0007) (0.0007) (0.0007)
Teams meeting total session time 0.0109∗∗ 0.0108∗∗ 0.0107∗∗ 0.0127∗∗
(0.0030) (0.0030) (0.0032) (0.0030)
Teams meetings attended 0.0068∗∗ 0.0067∗∗ 0.0040 0.0050∗∗
(0.0018) (0.0018) (0.0020) (0.0018)
Word total session time 0.0099 0.0107 0.0132 0.0213∗∗
(0.0068) (0.0068) (0.0073) (0.0070)
Others’ documents read/edited before completed 0.0590∗∗ 0.0574∗∗ 0.0340∗∗ 0.0570∗∗
(0.0121) (0.0121) (0.0123) (0.0121)
Number of email recipients -0.0009 -0.0009 -0.0008 -0.0009
(0.0004) (0.0004) (0.0003) (0.0004)
Number of unique people met 1-1 (total) -0.0004 -0.0004 0.0003 -0.0004
(0.0004) (0.0004) (0.0004) (0.0004)
Share of coworkers with Copilot license 0.0439 0.0528 0.0988∗∗ 0.0384
(0.0239) (0.0235) (0.0226) (0.0237)
Fixed-effects
Firm Yes Yes Yes
Industry Yes Yes
Fit statistics
Observations 2,718 2,718 2,718 3,167 3,167 3,167 2,718
R2 0.09294 0.09412 0.23525 0.18860 0.18388 0.05045 0.13906
Within R2 0.05276 0.00578 0.09072
Significance: ** $p<0.01$, * $p<0.05$
Note: This table shows OLS regressions results using individuals’ pre-experiment behaviors, share of coworkers with a Copilot license, and firm and industry fixed effects to predict mean post-period Copilot weekly usage among treated workers. All variables are averages over the pre-period other than “number of unique people met 1-1,” which is a total over a fixed pre-period.
Table A2: Predicting Weekly Copilot Use
Emails read -0.0003∗∗ -0.0003∗∗ -0.0002∗∗ -0.0004∗∗
($5.5\times 10^{-5}$) ($5.53\times 10^{-5}$) ($5.75\times 10^{-5}$) ($5.61\times 10^{-5}$)
Unique email threads replied to -0.0013∗∗ -0.0014∗∗ -0.0010 -0.0009
(0.0005) (0.0005) (0.0005) (0.0005)
Teams meeting total session time 0.0097∗∗ 0.0096∗∗ 0.0105∗∗ 0.0121∗∗
(0.0021) (0.0021) (0.0023) (0.0022)
Teams meetings attended 0.0074∗∗ 0.0074∗∗ 0.0041∗∗ 0.0053∗∗
(0.0013) (0.0013) (0.0014) (0.0013)
Word total session time 0.0127∗∗ 0.0131∗∗ 0.0134 0.0218∗∗
(0.0047) (0.0048) (0.0051) (0.0049)
Others’ documents read/edited before completed 0.0498∗∗ 0.0491∗∗ 0.0341∗∗ 0.0534∗∗
(0.0084) (0.0084) (0.0087) (0.0084)
Number of email recipients -0.0007∗∗ -0.0007∗∗ -0.0008∗∗ -0.0008∗∗
(0.0002) (0.0002) (0.0002) (0.0002)
Number of unique people met 1-1 (total) -0.0002 -0.0002 0.0002 -0.0003
(0.0003) (0.0003) (0.0003) (0.0003)
Share of coworkers with Copilot license 0.0236 0.0546∗∗ 0.1008∗∗ 0.0361
(0.0169) (0.0172) (0.0164) (0.0169)
Fixed-effects
Calendar week Yes Yes Yes Yes Yes Yes Yes Yes
Event month Yes Yes Yes Yes Yes Yes Yes Yes
Firm Yes Yes Yes
Industry Yes Yes
Fit statistics
Observations 61,826 61,826 61,826 72,412 72,412 72,412 72,412 61,826
R2 0.05026 0.05038 0.09285 0.07779 0.01986 0.07616 0.03432 0.06365
Within R2 0.03130 0.03141 0.01595 0.00177 0.02973
Significance: ** $p<0.01$, * $p<0.05$
Note: This table shows OLS regressions results using individual pre-experiment behaviors, share of coworkers with a Copilot license, firm industry fixed effects, and calendar-week and event-month fixed effects to predict post-period Copilot weekly usage among treated workers. All variables are averages other than “number of unique people met 1-1,” which is a total over a fixed pre-period.
C.2 Outcomes
Similar to Figure 2, Figure A4 shows event studies for the other variables in Table 2.
(a) Concentration Time
(b) Outlook Out-of-hours Time
(c) Email Threads Replied To
(d) Time to Reply
(e) %Meetings Early/Late
(f) Recurring Meeting Time
(g) Document TTC
(h) Collaborative Doc. TTC
(i) Non-Collab. Doc. TTC
(j) Completed Documents
(k) Completed Collab. Doc.
(l) Others’ Doc. Read/Edited
Figure A4: Event Study Plots for the Effect of Copilot on Other Outcomes
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table 2.
Table A3 shows the same analysis as Table 2 for some additional outcome variables. The decrease in emails read individually is unsurprising if Copilot is summarizing email threads (or all new emails) for people. Total emails replied to is another measure of reply activity where while the difference is statistically significiant, it is practically very small ($<2\%$).
We also look at a bunch of different types of meetings. While we do see some increases for big meetings, the event studies in Figure LABEL:fig:app_event_studies2 show that these increases only happened at the very start of the study period and seem likely due to the Copilot trainings that many companies did.
We also ran a balance test of workers’ pre-period behavior between treatment and control groups and found no significant differences across all outcomes.
Table A3: Effect of Copilot
Metric Pre-period mean (SD) OLS (SE) IV (SE) Workers $\mathbf{n}$
Emails read 155.57 -8.77** -20.07** 5,788 253,412
(99.60) (0.76) (1.75)
Total emails in email threads replied to 3.79 -0.036** -0.073** 6,111 285,332
(0.73) (0.010) (0.021)
[…]
Total Teams meetings attended 9.92 0.11 0.23 5,698 299,922
(6.14) (0.059) (0.12)
Total recurring meetings 4.14 0.046 0.091 5,698 254,319
(3.01) (0.026) (0.052)
Big short meetings 1.04 0.070** 0.14** 5,698 254,319
(1.02) (0.011) (0.021)
Big long meetings 0.87 0.035** 0.069** 5,698 254,319
(0.68) (0.008) (0.016)
Small short meetings 6.81 0.038 0.074 5,698 254,319
(4.23) (0.039) (0.077)
Small long meetings 1.15 -0.004 -0.007 5,698 254,319
(0.80) (0.010) (0.019)
[…]
Others’ documents read/edited 0.44 0.030 0.053 2,150 91,213
(0.64) (0.014) (0.025)
Significance: ** $q<0.01$, * $q<0.05$
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2) and the LATE coefficients (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$.
(a) Emails Read
(b) Emails Replied to
(c) Teams Meetings
(d) Big Short Meetings
(e) Big Long Meetings
(f) Recurring Meetings
(g) Small Short Meetings
(h) Small Long Meetings
Figure A5: Event Study Plots for the Effect of Copilot on Table A3 Outcomes
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table A3.
Table A4: Pre-period Balance Test
Metric Control Mean Treated Mean $q$-value
Outlook total session time 11.59 11.48 1
Total concentration time 24.55 24.67 1
Outlook out-of-hours session time 2.11 2.05 1
Unique email threads replied to 13.07 13.36 1
Time to reply from email delivery 14.63 14.89 1
Teams meeting total time 5.16 5.26 1
% Teams meetings attended late/early 0.29 0.30 1
Recurring Teams meetings time 2.28 2.35 1
Word total session time 1.37 1.33 1
Avg. document TTC 181.45 170.02 1
Avg. collaborative document TTC 270.14 286.12 1
Avg. non-collaborative document TTC 149.45 138.37 1
Completed documents 0.79 0.76 1
Emails read 156.26 154.92 1
Total emails in email threads replied to 3.79 3.79 1
Total Teams meetings attended 9.83 10.00 1
Total recurring meetings 4.07 4.21 1
Big short meetings 1.03 1.06 1
Big long meetings 0.86 0.88 1
Small short meetings 6.72 6.89 1
Small long meetings 1.14 1.17 1
Others’ documents read/edited 0.44 0.43 1
Number of email recipients 4.08 4.59 1
Number of unique people met 1-1 (total) 15.25 15.53 1
Note: This table shows average pre-period mean weekly outcomes for control and treatment group workers. No significant differences were found from a t-test on individual workers’ pre-period means. All time metrics in hours. A worker is a primary editor if they make the most edits on a document. Time to complete (TTC) and completed document counts are calculated only for documents where the target worker is the primary editor.
C.3 Splitting Sample by Firm Adoption
Table A5 is similar to Table 3 in the paper, but instead of splitting individuals by the share of their coworkers who had a Copilot licenses, it splits by whether the average Copilot usage rate at the firm was above or below the median of 45%. As one would expect the OLS coefficients are smaller at the firms where people used Copilot less – less usage leads to a smaller effect.
The IV results in Table A5 are somewhat hard to interpret because they are basically rescaling the ITT (OLS) results by the share of workers using Copilot while simultaneously splitting on that share. As shown in Figure 1(b), at low-usage firms a larger share of those using Copilot are using Outlook Copilot, which probably explains the larger IV coefficients for Outlook variables. It is not that Copilot usage actually had a larger effect at the lower-usage firms, but we are in some sense over-measuring the relevant (i.e. Outlook) Copilot usage at the high-usage firms.
Table A5: Effect of Copilot - above and below median firm usage of Copilot
Metric Pre-period OLS (SE) IV (SE) $\mathbf{n}$
mean (SD)
Above Below Above Below Above Below Above Below
Outlook total session time 11.32 (6.21) 11.80 (6.60) -1.47** (0.096) -1.21** (0.099) -2.46** (0.16) -4.22** (0.35) 156,648 124,635
Total concentration time 24.72 (5.97) 24.47 (6.45) 0.94** (0.083) 0.86** (0.089) 1.32** (0.14) 3.19** (0.33) 137,014 110,893
Outlook out-of-hours session time 2.00 (1.59) 2.19 (1.71) -0.34** (0.026) -0.25** (0.027) -0.56** (0.043) -0.88** (0.094) 156,648 124,635
Unique email threads replied to 13.12 (10.87) 13.34 (12.11) -0.31* (0.11) -0.13 (0.12) -0.50* (0.18) -0.41 (0.37) 181,898 143,574
Time to reply from email delivery 14.56 (4.50) 15.03 (5.18) -0.38** (0.12) -0.21 (0.14) -0.61** (0.19) -0.63 (0.43) 160,590 124,742
[…]
Teams meeting time 5.56 (3.23) 4.73 (3.29) 0.030 (0.045) 0.056 (0.050) 0.048 (0.072) 0.17 (0.15) 122,568 86,186
% Teams meetings attended late/early 0.29 (0.097) 0.30 (0.12) 0.010** (0.002) 0.002 (0.003) 0.015** (0.004) 0.005 (0.009) 107,111 71,107
Recurring Teams meetings time 2.53 (1.72) 2.02 (1.60) 0.018 (0.021) 0.006 (0.023) 0.028 (0.033) 0.019 (0.068) 148,836 105,483
[…]
Word total session time 1.42 (0.96) 1.24 (0.87) 0.044 (0.030) -0.025 (0.034) 0.066 (0.045) -0.068 (0.093) 71,076 43,787
Avg. document TTC 176.35 (84.33) 174.26 (80.63) -9.05 (6.67) -3.13 (9.04) -12.85 (9.46) -7.72 (22.25) 22,877 13,646
Avg. collaborative document TTC 282.10 (115.79) 271.15 (111.07) -37.51 (16.59) -30.61 (26.14) -52.44 (23.14) -65.23 (55.66) 7,429 3,604
Avg. non-collaborative document TTC 141.39 (80.15) 147.65 (74.08) -5.77 (6.51) -0.51 (8.63) -8.18 (9.23) -1.27 (21.33) 19,111 11,864
Completed documents 0.78 (0.73) 0.77 (0.77) 0.047* (0.021) -0.003 (0.028) 0.070* (0.032) -0.008 (0.076) 56,358 34,855
Completed collaborative documents 0.16 (0.25) 0.12 (0.18) 0.007 (0.007) 0.017 (0.008) 0.011 (0.011) 0.045 (0.023) 56,358 34,855
Significance: ** $q<0.01$, * $q<0.05$
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Columns 1 & 2), the intent-to-treat coefficients for the effect of Copilot (Columns 3 & 4), and the LATE estimates (Columns 5 & 6) separately for workers at firms who had above or below median Copilot usage in the post-period. All time metrics in hours. A worker is a primary editor if they make the most edits on a document. Time to complete (TTC) and completed document counts are calculated only for documents where the target worker is the primary editor.